If A is a unitary matrix, then all its eigen values are 1.

True or false with full explanation


1
Expert's answer
2021-02-24T07:07:16-0500

If A is a unitary matrix, then all its Eigen values are 1.

this statement is true because:

The eigenvalues and eigenvectors of unitary matrices have some special properties.

If U is unitary, then "U\nU\n\u2020\n=\nI"  Thus, if

"U\n|\nv\n\u27e9\n=\n\u03bb\n|\nv\n\u27e9"  (1)

then also

"\u27e8\nv\n|\nU\n\u2020\n=\n\u27e8\nv\n|\n\u03bb\n\u2217\n." (2)

Combining (1) and (2) leads to

"\u27e8\nv\n|\nv\n\u27e9\n=\n\u27e8\nv\n|\nU\n\u2020\nU\n|\nv\n\u27e9\n=\n\u27e8\nv\n|\n\u03bb\n\u2217\n\u03bb\n|\nv\n\u27e9\n=\n|\n\u03bb\n|\n2\n\u27e8\nv\n|\nv\n\u27e9"

Assuming "\u03bb\n\u2260\n0", we thus have

"|\n\u03bb\n|\n^2\n=\n1\n."

Thus, the eigenvalues of a unitary matrix are unimodular, that is, they have norm 1, and hence can be written as "e^{i\alpha}" for some "\alpha"